could someone explain to me how to find the parts of a rectangle step by step this is the problem:
The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., find the width of the rectangle.

Answer by rapaljer(4667)

The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., find the width of the rectangle.

Let x = width of the rectangle (since this is at the end of the first sentence!)
2x+3 = length of the rectangle

2W + 2L = Perimeter
2(x) + 2(2x+3) = 36
2x+4x+6 = 36
6x +6=36
6x=30
x=5 width
2x+3=Length
2(5) + 3= 10+3 = 13 Length

Check:
2W + 2L = P
2(5) + 2(13) = 10+26 = 36
It checks!

For additional explanation on word problems, especially rectangle problems, please see my Lesson Plan in algebra.com, or go to my website by clicking on my tutor name "rapaljer" anywhere in algebra.com, then look for MATH IN LIVING COLOR, look in Basic Algebra, Chapter 1. I hope what I have on my website will be helpful to you on this topic and many others about a variety of math concepts.

R^2 at SCC